# Provably Efficient Online RLHF with One-Pass Reward Modeling

### v2

@@ 4.3 Theoretical Guarantee
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big{\{}\big{\langle}g_{t}(\widetilde{\theta}_{t}),\theta\big{\rangle}+\frac{1}{\eta}\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})\Big{\}},$
where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is Bregman divergence. Based on the analysis of online mirror descent, we have the following lemma.
@@
Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define $\mathcal{C}_{t}=\{\theta\in\Theta\mid\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big{(}\sqrt{d}(\log(t/\delta))^{2}\big{)}\}$. Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}\right]\geqslant 1-\delta$.
Comparison with MLE.  For the MLE estimator in Eq. (3), prior works (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024) have shown $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\widetilde{\mathcal{O}}(\kappa\sqrt{d})$, where $V_{t}=\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$. By the definition of $\mathcal{H}_{t}$, it holds that $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Lemma 1 implies $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\sqrt{\kappa}\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}(\sqrt{\kappa d})$. This result shows that Lemma 1 improves upon previous bounds by at least a factor of $\sqrt{\kappa}$.
@@ 5 Applications in Three Online RLHF Scenarios
(a) Passive Data Collection
(b) Active Data Collection
(c) Deployment-Time Adaptation
[…]
@@ 5.1 Online RLHF with Passive Data Collection
where $\rho$ is the context distribution. The policy $\pi_{T+1}$ is selected as $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$. The detailed procedure is present in Algorithm 2, and we show it enjoys the following guarantee:
@@
The term $\big{\lVert}\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\big{\rVert}_{\mathcal{H}_{T+1}^{-1}}$ is referred to “concentrability coefficient” in the literature. It measures the distribution shift between the optimal policy and the collected data.
@@
Our result improves upon Zhu et al. (2023) in both statistical and computational efficiency. Specifically, since $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Theorem 2 directly implies a $\widetilde{\mathcal{O}}(\sqrt{d\kappa}\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ guarantee, improving their $\widetilde{\mathcal{O}}(\sqrt{d}\kappa\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ result by a factor of $\sqrt{\kappa}$. Regarding computational efficiency, their algorithm has a total storage complexity of $\mathcal{O}(T)$ and a time complexity of $\mathcal{O}(T\log T)$, leading to an amortized per-iteration cost of $\mathcal{O}(\log T)$. In contrast, our algorithm maintains a strict $\mathcal{O}(1)$ complexity per iteration, offering a substantial computational advantage.
Algorithm 3 Active Data Collection
0:  Regularization $\lambda$, step size $\eta$
1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\mathcal{H}_{1}=\lambda I$
2:  for $t=1,2,\ldots,T$ do
3:     Choose $(x_{t},a_{t},a_{t}^{\prime})$ as Eq. (7), observe $y_{t}$
[…]
5:  end for
6:  $\pi_{T+1}(x)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\widetilde{r}_{T+1}(x,a)$
Algorithm 4 Deployment-Time Adaptation
0:  Regularization $\lambda$, step size $\eta$
1:  Initialize $\widetilde{\theta}_{1}=0$ and $\mathcal{H}_{1}=\lambda I$.
2:  for $t=1,2,\ldots,T$ do
3:     Observes the context $x_{t}$.
4:     Selects $a_{t}$ and $a_{t}^{\prime}$ as Eq. (9) and Eq. (10)
5:     Observe the preference feedback $y_{t}$
[…]
7:  end for
@@ 5.2 Online RLHF with Active Data Collection
Active Data Collection.  At each iteration, we obtain a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$. We then update the reward model using our one-pass reward modeling method as in Eq. (5). To guide data acquisition, we actively select the query with the highest uncertainty under the current reward model. Specifically, we select the next query by solving:
@@ 5.3 Online RLHF with Deployment-Time Adaptation
In this section, we consider the deployment-time adaptation setting, where users provide input contexts in an online manner, and the learner generates responses while simultaneously collecting feedback to improve the model. In this scenario, the learner faces a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that yield informative feedback to facilitate continual model improvement. To this end, we choose the cumulative regret as the performance measure, defined as
${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{1}{2}\left(r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)\right)\right),$ (8)
where $\pi^{*}$ is the optimal policy. This measure differs from the one introduced by Ji et al. (2024), which evaluates the gap between th e optimal action and a single selected action. Our measure is designed to ensure that both actions are sufficiently effective, thereby delivering a high-quality user experience. While better aligns with deployment needs, it also poses greater challenges. Specifically, to optimize the measure in Ji et al. (2024), the learner can exploit one high-reward action while using the other for uniform exploration. In contrast, our setting imposes a stricter requirement: both actions must be good, thereby limiting the ability to sacrifice one action purely for exploration.
Action selection.  At each iteration, given a new preference data $(x_{t},a_{t},a_{t}^{\prime},y_{t})$, the learner updates the reward model using our one-pass reward modeling method as in Eq. (5). The learner must select queries that are both informative and with high rewards. To address this, we choose the first action $a_{t+1}$ to maximize the estimated reward based on the estimator $\widetilde{\theta}_{t+1}$, i.e.,
@@
For any $\delta\in(0,1]$, set parameters as in Lemma 1, Algorithm 4 ensures with probability at least $1-\delta$, the regret satisfies
${\operatorname*{Reg}}_{T}\leq\widetilde{\mathcal{O}}\big{(}d\sqrt{{\kappa}{T}}\big{)}.$
@@ 6 Practical Implementation
In this section, we introduce the practical implementation of our proposed algorithm.
@@ 6.1 Computation of Inverse Hessian
The OMD update in Eq. (5) requires computation of the inverse of the Hessian matrix. Omitting the projection operation, Eq. (5) can be rewritten as $\widetilde{\theta}_{t+1}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$. Computing the full Hessian inverse ~$\widetilde{\mathcal{H}}_{t}^{-1}$ directly incurs a time complexity of $\mathcal{O}(d^{3})$, which is prohibitive for LLMs.
This cost can be reduced to $\mathcal{O}(d^{2})$ by applying the Sherman-Morrison-Woodbury formula, leveraging the fact that the Hessian is a rank-one update. Specifically, for a matrix of the form $A+\mathbf{x}\mathbf{x}^{\top}$ where $A$ is invertible and $\mathbf{x}$ is a vector, the inverse is given by $(A+\mathbf{x}\mathbf{x}^{\top})^{-1}=A^{-1}-\frac{A^{-1}\mathbf{x}\mathbf{x}^{\top}A^{-1}}{1+\mathbf{x}^{\top}A^{-1}\mathbf{x}}$, requiring only $\mathcal{O}(d^{2})$ time. Nevertheless, even this reduced complexity is costly for large models.
To further reduce the computational burden to $\mathcal{O}(d)$, we employ the Hessian-vector product technique combined with conjugate gradient descent (Boyd and Vandenberghe, 2004). Instead of explicitly computing ~$\widetilde{\mathcal{H}}_{t}^{-1}$, we solve the linear system ~$\widetilde{\mathcal{H}}_{t}v=g_{t}(\widetilde{\theta}_{t})$ iteratively, where $v$ is the solution we seek. The HVP operation ~$\widetilde{\mathcal{H}}_{t}v$ for any vector $v$ can be computed efficiently using twice differentiation: ~$\widetilde{\mathcal{H}}_{t}v=\nabla_{\theta}\big{(}\nabla_{\theta}\mathcal{L}_{t}(\theta)^{\top}v\big{)}\big{|}_{\theta=\widetilde{\theta}_{t}}+\lambda_{t}v$, where $\mathcal{L}_{t}(\theta)$ is the function whose Hessian matrix at $\widetilde{\theta}_{t}$ is ~$\widetilde{\mathcal{H}}_{t}$. By leveraging this efficient computation, the conjugate gradient method iteratively refines the solution. In practice, finding $\mathcal{L}_{t}(\theta)$ is challenging as ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ involve complex structure. A key insight is ~$\widetilde{\mathcal{H}}_{t}$ can be viewed as the cumulative Hessian of the $\ell_{t}(\theta)$ approximately. This motivates the use of a more tractable surrogate $\ell_{t}(\theta)$ in Eq. (3) while keeping the update. To further approximate the effect of curvature accumulation, we adopt an adaptive damping scheme that approximately preserves historical Hessian information while incorporating new Hessian updates: $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a monotonic increasing function, such as linear functions.
@@ 6.2 Computation of Model Uncertainty
In both online RLHF with active data collection and deployment-time adaptation, our algorithm utilizes uncertainty-driven query selection strategies. While quantifying uncertainty using the local norm induced by the Hessian matrix offers strong theoretical guarantees, it is computationally prohibitive in practice. To address this challenge, we adopt a rejection sampling-based approximation, a technique commonly employed for exploration in the literature (Nakano et al., 2021; Gulcehre et al., 2023; Dong et al., 2023, 2024). Specifically, given a prompt, we sample $n$ independent responses by the current model, then use the trained reward model to rank the responses. Then, we use different strategies to select the response for different settings. Specifically, In active data collection, the key insight is to identify and query samples that exhibit the greatest diversity in prompt action features. To this end, we select the response with the highest predicted reward and the response with the lowest predicted reward. In deployment-time adaptation, the core idea is to select the first action to maximize the estimated reward, while the second is chosen to balance high reward with sufficient divergence from the first. Concretely, we select the response with the highest predicted reward and another from the top-$1/q$ percentile of the reward, where $q$ is a hyperparameter.
@@ 7 Experiments
In this section, we empirically evaluate the performance of our proposed method.  [^1] Equal contribution. We first describe the experimental setup, and then present the empirical results.
7.1 Experiment setup
In our experiments, we employ the Llama-3-8B-Instruct 222huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and Qwen2.5-7B-Instruct 333huggingface.co/Qwen/Qwen2.5-7B-Instruct as the foundation model for reward model. We extract features $\phi(x,a)$ using the last layer of the model, and the dimension is $d=4096$. We use two datasets for evaluation. The first one is Ultrafeedback-binarized dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, a pre-processed version of the original Ultrafeedback dataset (Cui et al., 2023), a widely used benchmark for RLHF. It collects about $64,000$ prompts from diverse resources, including question answering, summarization, and dialogue generation. Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a preference label $y$. We also employ a mixed dataset, Mixture2 dataset 444huggingface.co/datasets/OpenRLHF/preference_dataset_mixture2_and_safe_pku, which combines a variety of preference datasets, including HH-RLHF, SHP, UltraFeedback, etc.
11footnotetext: The code is available at https://github.com/ZinYY/Online_RLHF
@@ 7.1 Experiment setup
Figure 2: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 3: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
(a) training loss
(b) training accuracy
(c) evaluation loss
(d) evaluation accuracy
Figure 4: For Llama-3-8B-Instruct on Mixture2 dataset with passive data collection, we report the comparison of MLE and our OMD-based method for the reward model about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
7.2 Experimental results
In this section, we present the experimental results, including passive data collection, active data collection and deployment-time adaptation.
7.2.1 Passive Data Collection
We randomly sample $T=30,000$ data points from the dataset for training the reward model. We evaluate both the standard MLE-based method, which uses a stochastic gradient descent (SGD) optimizer, and our proposed OMD-based method in terms of reward model loss and accuracy. Figures 23, and 4 present the results for three settings: (i) the Llama-3-8B-Instruct model on the Ultrafeedback-binarized dataset, (ii) the Qwen2.5-7B model on the same dataset, and (iii) the Llama-3-8B-Instruct model on the Mixture2 dataset. We report the results across four key metrics: (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy. Our method demonstrates faster convergence to a lower loss and achieves higher evaluation accuracy compared to the MLE baseline. These results highlight the superior statistical efficiency of the OMD-based approach, which achieves improved performance with fewer training samples.
@@ 7.2.1 Passive Data Collection
Figure 5: For Llama-3-8B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
(a) training loss
(b) evaluation accuracy
Method ACC (%) Time (s)
Rand-MLE 67.36 $\pm$ 0.5 4653 $\pm$ 121
Active-MLE 67.25 $\pm$ 0.4 4701 $\pm$ 103
Rand-OMD 66.80 $\pm$ 0.5 1312 $\pm$ 47 0
Ours 67.11 $\pm$ 0.5 1325 $\pm$ 54 0
(c) evaluation accuracy and training time
Figure 6: For Qwen2.5-7B-Instruct on Ultrafeedback-binarized dataset with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
7.2.2 Active Data Collection

### v3

@@ 4.3 Theoretical guarantee
$\widetilde{\theta}_{t+1}=\operatorname*{arg\,min}_{\theta\in\Theta}\Big\{\eta\big\langle g_{t}(\widetilde{\theta}_{t}),\theta\big\rangle+\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})\Big\},$
where $\psi_{t}(\theta)=\frac{1}{2}\lVert\theta\rVert_{\widetilde{\mathcal{H}}_{t}}^{2}$ is the regularizer and $\mathcal{D}_{\psi_{t}}(\theta,\widetilde{\theta}_{t})=\psi_{t}(\theta)-\psi_{t}(\widetilde{\theta}_{t})-\langle\nabla\psi_{t}(\widetilde{\theta}_{t}),\theta-\widetilde{\theta}_{t}\rangle$ is Bregman divergence. Leveraging the modern analysis of online mirror descent (Zhao et al., 2024; Zhang and Sugiyama, 2023), we derive the following estimation error bound.
@@
Let $\delta\in(0,1]$, set $\eta=(1/2)\log 2+\left(BL+1\right)$ and $\lambda=84\sqrt{2}\eta(dL^{2}+BL^{3})$, define $\mathcal{C}_{t}=\{\theta\in\Theta\mid\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\beta}_{t}\triangleq\mathcal{O}\big(\sqrt{d}\log(t/\delta)\big)\}$. Then, we have $\operatorname{Pr}\left[\forall t\geqslant 1,\theta^{*}\in\mathcal{C}_{t}\right]\geqslant 1-\delta$.
Comparison with MLE.  For the MLE estimator in Eq. (1), prior works (Zhu et al., 2023; Das et al., 2025; Ji et al., 2025) have shown $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\widetilde{\mathcal{O}}(\kappa\sqrt{d})$, where $V_{t}=\sum_{i=1}^{t-1}z_{i}z_{i}^{\top}+\lambda I$. By the definition of $\mathcal{H}_{t}$, it holds that $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Lemma 1 implies that $\|\theta-\widetilde{\theta}_{t}\|_{V_{t}}\leq\sqrt{\kappa}\|\theta-\widetilde{\theta}_{t}\|_{\mathcal{H}_{t}}\leq\widetilde{\mathcal{O}}(\sqrt{\kappa d})$. This result shows that Lemma 1 improves upon previous bounds by at least a factor of $\sqrt{\kappa}$.
@@ 5.1 Online RLHF with passive data collection
where $\rho$ is the context distribution. The policy $\pi_{T+1}$ is selected as $\pi_{T+1}=\operatorname*{arg\,max}_{\pi\in\Pi}\widetilde{J}_{T+1}(\pi)$. The detailed procedure is present in Algorithm 2, and we show it enjoys the following guarantee.
@@
The term $\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{\mathcal{H}_{T+1}^{-1}}$ is usually referred to “concentrability coefficient” in the literature. It measures the distribution shift between the optimal policy and the collected data.
@@
For statistical efficiency, since $\mathcal{H}_{t}\succeq\kappa^{-1}V_{t}$, Theorem 1 improves the $\widetilde{\mathcal{O}}(\sqrt{d}\kappa\cdot\lVert\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]\rVert_{V_{T+1}^{-1}})$ result of  Zhu et al. (2023) at least by a factor of $\sqrt{\kappa}$. Regarding computational efficiency, their algorithm has a total storage complexity of $\mathcal{O}(T)$ and a time complexity of $\mathcal{O}(T\log T)$, leading to an amortized per-iteration cost of $\mathcal{O}(\log T)$. In contrast, our algorithm maintains a strict $\mathcal{O}(1)$ complexity per iteration, offering a substantial computational advantage.
(a) Passive Data Collection
(b) Active Data Collection
(c) Deployment-Time Adaptation
Figure 1: Different settings of online RLHF. Contexts and actions selected by the environment () are shown in grey, while those selected by the algorithm () are highlighted in color.
@@ 5.2 Online RLHF with active data collection
Active Data Collection.  At each iteration, we select a triplet $(x_{t},a_{t},a^{\prime}_{t})$ to query for human feedback $y_{t}$, and then update the reward model using our one-pass method as defined in Eq. (3). To guide data acquisition, we adopt an active selection strategy that queries the sample with the highest predictive uncertainty under the current reward model. Specifically, the next query is chosen by solving:
@@ 5.3 Online RLHF with deployment-time adaptation
In this section, we consider the deployment-time adaptation setting, where users provide input contexts in an online manner, and the learner generates responses while simultaneously collecting feedback to improve the model. In this scenario, the learner faces a dual objective: selecting actions that maximize rewards to ensure a positive user experience, while also choosing actions that yield informative feedback to facilitate continual model improvement. To this end, we consider the measure: ${\operatorname*{Reg}}_{T}=\sum_{t=1}^{T}\left(r\left(x_{t},\pi^{*}(x_{t})\right)-\frac{1}{2}\left(r\left(x_{t},a_{t}\right)+r\left(x_{t},a^{\prime}_{t}\right)\right)\right)$, where $\pi^{*}$ is the optimal policy.
Action selection.  At each iteration, given a prompt $x_{t}$ from the user, the learner selects two actions $a_{t}$ and $a_{t}^{\prime}$ and obtain the feedback $y_{t}$. The learner must select actions that are both informative and with high rewards. To this end, we choose the first action $a_{t+1}$ to maximize the estimated reward, i.e.,
@@
For any $\delta\in(0,1]$, set parameters as in Lemma 1, Algorithm 4 ensures the regret satisfies ${\operatorname*{Reg}}_{T}\leq\widetilde{\mathcal{O}}\big(d\sqrt{{\kappa}{T}}\big)$ with probability at least $1-\delta$.
@@
Algorithm 3 Active Data Collection
0: Regularization parameter $\lambda$, step size $\eta$
1: Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\mathcal{H}_{1}=\lambda I$
2:for $t=1,2,\ldots,T$ do
3:  Choose $(x_{t},a_{t},a_{t}^{\prime})$ as Eq. (5), observe $y_{t}$
[…]
5:end for
[…]
Algorithm 4 Deployment-Time Adaptation
0: Regularization parameter $\lambda$, step size $\eta$
1: Initialize $\widetilde{\theta}_{1}=0$ and $\mathcal{H}_{1}=\lambda I$.
2:for $t=1,2,\ldots,T$ do
3:  Observes the context $x_{t}$.
4:  Selects $a_{t}$ and $a_{t}^{\prime}$ as Eq. (6) and Eq. (7)
5:  Observe the preference feedback $y_{t}$
[…]
7:end for
@@ 6 Practical Implementation
While the proposed one-pass algorithm completely removes the need to store historical data and achieves constant-time updates per iteration, its computational cost still exhibits an implicit dependence on the feature dimension $d$, which can become non-negligible in large-scale model optimization. To further alleviate this issue, we introduce in this section several empirical approximation techniques designed to reduce the effective dependence on dimensionality and enhance practical efficiency.
@@ 6.1 Computation of inverse Hessian
Although the OMD update in Eq. (3) enjoys one-pass property, it requires the computation of matrix inversion. Specifically, by omitting the projection operation, Eq. (3) can be rewritten as $\widetilde{\theta}_{t+1}=\widetilde{\theta}_{t}-\eta\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$ where $\widetilde{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\eta H_{t}(\widetilde{\theta}_{t})+\lambda I$. Computing the full $\widetilde{\mathcal{H}}_{t}^{-1}$ directly incurs a time complexity of $\mathcal{O}(d^{3})$, which is prohibitive for LLMs as $d$ is typically large.
This cost can be reduced to $\mathcal{O}(d^{2})$ by applying the Sherman-Morrison-Woodbury formula, leveraging the fact that the Hessian is a rank-one update. Specifically, for a matrix of the form $A+\mathbf{x}\mathbf{x}^{\top}$ where $A$ is invertible and $\mathbf{x}$ is a vector, the inverse is given by $(A+\mathbf{x}\mathbf{x}^{\top})^{-1}=A^{-1}-\frac{A^{-1}\mathbf{x}\mathbf{x}^{\top}A^{-1}}{1+\mathbf{x}^{\top}A^{-1}\mathbf{x}}$, requiring only $\mathcal{O}(d^{2})$ time. Nevertheless, even this reduced complexity remains costly for large models.
To further reduce the computational burden to $\mathcal{O}(d)$, we employ the Hessian-vector product technique combined with conjugate gradient descent (Boyd and Vandenberghe, 2004). Instead of explicitly computing $\widetilde{\mathcal{H}}_{t}^{-1}$, we define $v_{t}=\widetilde{\mathcal{H}}_{t}^{-1}g_{t}(\widetilde{\theta}_{t})$ and solve the linear system $\widetilde{\mathcal{H}}_{t}v_{t}=g_{t}(\widetilde{\theta}_{t})$ using the conjugate gradient method. The required matrix–vector product decomposes as $\widetilde{\mathcal{H}}_{t}v_{t}={\sum\nolimits_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})v_{t}}+{\lambda v_{t}}+{\eta H_{t}(\widetilde{\theta}_{t})v_{t}}$.
For the first term, materializing and storing all past Hessians $H_{i}(\widetilde{\theta}_{i+1})$ is infeasible. We therefore absorb their effect into the second term by replacing $\lambda$ with $\lambda_{t}=\lambda_{0}\cdot\min\{1,f(t/T)\}$, where $f(\cdot)$ is a monotonic increasing function, such as a linear or logarithmic function. The last term can be computed via the Pearlmutter trick as $H_{t}(\widetilde{\theta}_{t})v_{t}=\nabla_{\theta}\big(\nabla_{\theta}\ell_{t}(\theta)^{\top}v_{t}\big)\big|_{\theta=\widetilde{\theta}_{t}}$. Each iteration therefore requires only HVPs and vector operations, yielding an overall $\mathcal{O}(d)$ per-iteration cost with a small fixed number of iterations.
@@ 6.2 Computation of model uncertainty
In both online RLHF with active data collection and deployment-time adaptation, our algorithm utilizes uncertainty-driven query selection strategies. While quantifying uncertainty using the local norm induced by the inverse Hessian matrix offers strong theoretical guarantees, it is computationally prohibitive in practice. To address this challenge, we adopt a rejection sampling-based approximation, a technique commonly employed for exploration in the RLHF literature (Nakano et al., 2021; Gulcehre et al., 2023; Dong et al., 2023, 2024). Specifically, given a prompt, we sample $n$ independent responses by the current model, then use the trained reward function to rank the responses. Then, we use different strategies to select the response for different settings. Specifically, In active data collection, the key insight is to identify and query samples that exhibit the greatest diversity in prompt action features. To this end, we select the response with the highest predicted reward and the one with the lowest predicted reward. In deployment-time adaptation, the core idea is to select the first arm to maximize the estimated reward, while the second is chosen to balance high reward with sufficient divergence from the first. Concretely, we select the response with the highest predicted reward and another from the top-$1/q$ percentile of the reward to ensure diversity, where $q$ is a hyperparameter.
@@ 7 Experiments
Figure 2: For online RLHF with passive data collection, we report the comparison of MLE and our method about (a) training loss, (b) training accuracy, (c) evaluation loss and (d) evaluation accuracy.
@@
Figure 3: For online RLHF with active data collection, we report the comparison of different methods about (a) training loss, (b) evaluation accuracy and (c) final evaluation accuracy and training time.
In this section, we empirically evaluate the performance of our proposed method.  [^1] The code is available at https://github.com/ZinYY/Online_RLHF We first describe the experimental setup, and then present the empirical results.
7.1 Experiment setup
In our experiments, we employ the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct as the base model for reward model. We extract features $\phi(x,a)$ using the last layer of the model, and the dimension is $d=4096$. We use two datasets for evaluation. The first one is Ultrafeedback-binarized dataset, a pre-processed version of the original Ultrafeedback dataset (Cui et al., 2024), a widely used benchmark for RLHF. It collects about $64,000$ prompts from diverse resources, including question answering, summarization, and dialogue generation. Each data consists of a context $x$, two responses $a$ and $a^{\prime}$, and a preference label $y$. We also employ a mixed dataset, Mixture2 dataset (Dong et al., 2024), which combines a variety of preference datasets, including HH-RLHF, SHP, UltraFeedback, Capybara, etc. The dataset follows the same format as the UltraFeedback-binarized dataset.
7.2 Experimental results
We present the experimental results for Llama-3-8B-Instruct on the Ultrafeedback dataset. Due to page limits, more detailed results including comparisons with Adam, DPO, full model updates, additional models of Qwen2.5-7B-Instruct, and Mixture2 dataset are deferred to appendix.
Passive data collection.  We evaluate the performance of our proposed method in terms of the loss and accuracy of the reward model. We compare our OMD-based method with the MLE-based method. We randomly sample $T=30,000$ data points from the Ultrafeedback dataset for training. Figure 2 shows the loss and accuracy vs. the number of training samples. We observe that our method converges faster to a lower loss and achieves a higher evaluation accuracy compared to baselines. The improvement is particularly pronounced in the small-sample regime ($T<10,000$), where our method achieves a higher evaluation accuracy with the same amount of samples compared to MLE which employs conventional stochastic gradient descent (SGD) updates. This shows the superior statistical efficiency of our approach, achieving a better performance with fewer training samples.